OpenStudy (anonymous):

What is the wavelength (in nm) of radiation that has an energy content of 7.67 x 10^3 kJ/mol? I got myself confused on how to set it up.

OpenStudy (aaronq):

you have to get the value in energy terms (per photon, not per mole of photons)

OpenStudy (anonymous):

How do you set that up? Do you use the E = hv formula?

OpenStudy (aaronq):

yeah, but you need to divide by avogadro's number first

OpenStudy (anonymous):

For which part?

OpenStudy (anonymous):

\[\frac{ 7.67 \times 10^3 kJ }{ mol } \times \frac{ 1000J }{ 1kJ } \times \frac{ 1mol }{6.022 \times 10^{23}photons }\]. You would use this dimensional analysis to find energy per photon (J/photon). This would be E Next use E=hv h is a Planck constant Find v which is frequency Then use \[c=\lambda v\] where c is speed of light. It has a constant of 3.00 x 10^8m/s You are to find lambda which is wavelength. It would be in meters. 1m=1.0 x 10^9nm Good luck solving it.

OpenStudy (anonymous):

1.56 x 10^11. Would that be right?

OpenStudy (anonymous):

You mean 1.56 x 10^1 Because thats what I got.

OpenStudy (anonymous):

Yeah. Meant that.

OpenStudy (anonymous):

How do you tell where it is on the electromagnetic spectrum?

OpenStudy (anonymous):

Alright. Thank you.

OpenStudy (anonymous):

You write the answer in meters Then you would use this chart. http://www2.lbl.gov/MicroWorlds/ALSTool/EMSpec/EMSpec2.html

OpenStudy (anonymous):

No problem! It can be challenging at first. Good luck on your work.

OpenStudy (anonymous):

Thanks!